# Graph-Based Approximate Nearest Neighbor Search Revisited: Theoretical Analysis and Optimization

### v2

@@ 3 Degree Bounds of SNG
As outlined at the beginning of this section, our goal is to show that the total number of iterations $T$, which corresponds to the degree, satisfies $T=O(n^{2/3+\epsilon})$. We divide the SNG construction into two parts, based on the number of processed points $S^{\prime}_{t}$: (1) When $|S_{t}|<n-n^{1-\nu}$, for some $\nu\in(2/3,1)$, Lemma 8 ensures that the number of iterations to reach this stage is $t_{1}=O(n^{\nu})$ with probability 1. (2) When $|S^{\prime}_{t}|\geq n-n^{1-\nu}$, the process is within the plateau phase. Applying DEM, the discrete process $S^{\prime}_{t}$ can be approximated by the solution to an ODE. Solving this ODE, the number of iterations required to reach n1n-1 processed points in this phase is $O(n^{1-\nu})$. Adding both phases gives $T=\ O(n^{\nu})+O(n^{1-\nu})=O(n^{\nu})$, yielding the bound $O(n^{2/3+\epsilon})$ for any small $\epsilon>0$. ∎
4 Search Bounds
We now turn to the analysis of search on the constructed SNG. Our goal is to show that with high probability, GreedySearch converges to an approximate nearest neighbor in $O(\log n)$ steps. The full proof of Theorem 2 is provided in Appendix C.5.
@@ 4 Search Bounds
Given a dataset $P$ with $n$ points in $d$-dimensional space, with high probability $1$, the SNG search converges in $O(\log n)$ steps.
@@
To analyze the expected path length $k$, we analyze the distribution of neighbors around the destination $v_{k}$, assuming they are uniformly distributed in a $d$-dimensional ball of radius $\rho_{0}$. We partition this ball into concentric annular layers and let $\eta_{i}$ denote the number of neighbors lying in the $i$-th layer. Each $\eta_{i}$ follows a binomial distribution proportional to the shell volume, and the total number of neighbors is bounded by $\eta=O(n^{2/3+\epsilon})$ from Theorem 1.
@@
Using volume arguments, we upper bound the expected total number of neighbors located within the ball of radius $\|v_{k}-v_{0}|$ by the expected number of dataset points that fall into the same region. Specifically, we show $E[k\cdot\eta]\leq\frac{n(n+1)^{(d-1)/d}\rho_{0}^{d+1}\cdot E\left[\left((\rho_{0}-\triangle r)/\rho_{0}\right)^{d(k-1)}\right]}{\triangle r^{d+1}\cdot d},$ where $\triangle r$ is a lower bound on per-step radial improvement. Applying Jensen’s inequality, we define an auxiliary function $g(x)$ whose root bounds $E[k]$. We demonstrate $g(k)<0$, note that $g(x)$ is monotonically increasing, and $g(\log n)>0$, which together imply that $k<\log n$, hence $E[k]=O(\log n)$. ∎
5 Truncation Parameter Optimization
In the preceding sections, we establish theoretical guarantees for the SNG structure. A practical and widely adopted realization of SNG is implemented in the Vamana algorithm, which serves as the core indexing component in DiskANN system NEURIPS2019 . In the following, we focus on optimizing the truncation parameter specifically for Vamana algorithm (see Appendix E for algorithm pseudo code).
5.1 Complexity Analysis
Total Complexity of Vamana. The following are the graph construction steps of the Vamana algorithm. It begins by constructing an initial randomized $R$-regular graph with truncation parameter $R$. For each data point $p$, GreedySearch is performed from the dataset centroid to gather a candidate set for subsequent pruning. Following candidate collection, the graph is refined through iterative SNG pruning, truncation limits the number of pruning iterations to $R$. Additionally, reverse edges are added to neighbors of each point to improve search connectivity, which may cause degree overruns. Such overruns trigger further pruning. Detailed complexity analyses for GreedySearch and pruning are deferred to Appendix F, the overall construction complexity of the Vamana algorithm is expressed as:
$n(C_{1}\cdot R\log n+b_{1}+C_{2}(R\cdot R\log n/\alpha)+b_{2}+C_{3}(\alpha\cdot R\cdot R^{2})+b_{3}).$
where $C_{i}$$b_{i}$ are implementation-dependent constants.
5.2 Optimization Principle
As discussed in Section 2.1, the pruning parameter $\alpha$ is typically predefined. In contrast, the truncation parameter $R$ introduces an intrinsic trade-off: increasing $R$ improves search accuracy but also increases construction complexity. The optimal balance between these objectives varies across applications, depending on the tolerance of each application. This reflects the absence of a universal standard trade-off.
Interestingly, our empirical observations reveal that the construction complexity exhibits a non-monotonic relationship with respect to $\alpha$: it initially decreases and then increases. This suggests the existence of an optimal $\alpha$ that minimizes construction cost for a given $R$. Motivated by this, we propose computing the value of $R$ that achieves optimal efficiency with respect to a target $\alpha$, leading to improvements beyond the naive trade-off. We refer to this as the marginal optimality principle.
Algorithm 1 Optimization of Truncation Parameter $R$
1:Input: Test pruning parameter $\alpha_{1}$, target pruning parameter $\alpha_{2}$, number of data points $n$
2:Output: Optimized truncation parameter $R^{*}$
3: Initialize $R\leftarrow n^{2/3}$
@@ 5.2 Optimization Principle
5: Compute average degree $\bar{R}$ from the resulting graph
6: Compute scaling constant $K’\leftarrow\frac{\alpha_{1}^{2}\cdot\bar{R}}{\log n}$
7: Compute optimized truncation $R^{*}\leftarrow\frac{K^{\prime}\cdot\log n}{\alpha_{2}^{2}}$
8:Return $R^{*}$
@@
Our theoretical analysis, by setting the derivative of the construction complexity with respect to $\alpha$ to zero, yields the relationship $R^{*}\propto\log n/\alpha^{2}$. This approach is inspired by the complexity-aware tuning strategies used in prior work, such as TKDE . The goal of our algorithm is to determine the constant of proportionality. To this end, we begin with an initial construction using $R=n^{2/3}$ under a test pruning parameter $\alpha_{1}$. This choice corresponds to a known theoretical upper bound and approximates the structure of a non-truncated graph. We then measure the average degree $\bar{R}$ of the constructed graph to estimate the scaling coefficient $K^{\prime}$ in the formula of $R^{*}$. Having obtained this coefficient, we directly compute the optimal truncation parameter $R^{*}$ for the target pruning parameter $\alpha_{2}$.
6 Experiment
Experiment Setup. We evaluate our parameter optimization method against the traditional binary search baseline through two sets of experiments: (1) End-to-end parameter selection and graph construction, where we measure the total index construction time; and (2) Query-time performance evaluation, where we assess the resulting indices using both latency and Recall@@10 metrics [^6] Recall@@10 measures the proportion of queries for which at least one true nearest neighbor appears among the top-10 retrieved candidates. Formally, Recall@$k$ = $\frac{|S\cap T|}{|T|}$, where $S$ is the retrieved set and $T$ is the ground truth.. All the experiments are conducted on a machine with an Intel i7-14700KF CPU (20 cores), 64GB DDR4 RAM, and a Samsung 990 EVO Plus 2TB SSD. Codes are available at https://anonymous.4open.science/r/Graph-based-ANNS-F168.
We use Vamana NEURIPS2019 as the baseline—a generalization of Sparse Neighborhood Graphs with a tunable pruning parameter $\alpha$—due to its strong empirical performance on large-scale datasets. All experiments are conducted on three standard public benchmarks: SIFT1M, GIST1M DBLP:journals/pami/JegouDS11 , and DEEP1M DEEP1M . Dataset statistics are summarized in Table 2.
To further assess the distributional robustness of our theoretical results, we conduct an experiment on a synthetic dataset generated from a Gaussian Mixture Model (GMM). GMMs are known to approximate a broad class of real-world distributions effectively GMM_ref . We create a dataset with 100,000 points in 8 dimensions, construct an SNG on it, and analyze the resulting degree distribution. The goal is to empirically verify whether the sparsity property and degree bound established under uniform distributions (Theorem 1) is scalable for non-uniform distributions.
6.1 Index Construction Time Overhead
In the truncated version of the SNG, graph construction consists of two phases: parameter selection and index building. We compare our analytical parameter optimization method with the traditional binary search approach by separately measuring the parameter selection time (denoted as tuning time) and the graph construction time using the Vamana algorithm (denoted as construction time) across various values of the pruning parameter $\alpha$. As shown in Figure 1, our method consistently accelerates the overall construction pipeline. Specifically, it reduces the total construction time by over 47% in all tested settings, with the majority of the gain attributed to the significant reduction in tuning time. In particular, the tuning time is reduced by more than 50% , demonstrating the efficiency of our theoretical optimization approach.
Figure 1: Comparison of total graph construction time
6.2 Search Performance
We compare the optimized truncation parameter $R$, derived from our analytical method, against the baseline values obtained via binary search across three benchmark datasets: SIFT1M, GIST1M, and DEEP1M. Evaluation is conducted using recall and query latency. The results are presented in Figures 2, 3, and 4, respectively. Across all datasets, our method consistently achieves higher recall at comparable latency levels. For example, on GIST1M, our method maintains similar recall under $\alpha=1.00$ and $\alpha=1.25$, and achieves a 1.52% absolute improvement in recall at 228 ms latency under $\alpha=1.50$. On DEEP1M, under $\alpha=1.00$, our approach improves recall by at least 1.9% at just 5 ms latency. These results demonstrate that our parameter optimization method matches or outperforms binary search-based heuristics across datasets with varying characteristics, while being significantly more efficient in tuning.
@@ 6.2 Search Performance
Figure 2: Query performance across varying $\alpha$ values on SIFT1M

### v3

@@ 4.4. Main Degree Bound
As outlined at the beginning of this section, our goal is to show that the total number of iterations $T$, which corresponds to the degree, satisfies $T=O(n^{2/3+\epsilon})$. We divide the SNG construction into two parts, based on the number of processed points $|S^{\prime}_{t}|$:
(1) When $|S_{t}|<n-n^{1-\nu}$, for some $\nu\in(2/3,1)$, Lemma 3 ensures that the number of iterations to reach this stage is $t_{1}=O(n^{\nu})$ with probability one.
(2) When $|S^{\prime}_{t}|\geq n-n^{1-\nu}$, the process is within the plateau phase. Applying DEM, the discrete process $S^{\prime}_{t}$ can be approximated by the solution to an ODE. Solving this ODE, the number of iterations required to reach n1n-1 processed points in this phase is $O(n^{1-\nu})$.
Adding both phases gives
(19) $T=\ O(n^{\nu})+O(n^{1-\nu})=O(n^{\nu}),$
yielding the bound $O(n^{2/3+\epsilon})$ for arbitrarily small $\epsilon>0$. ∎
5. Search Bounds
We now turn to the analysis of query processing on the constructed SNG. Our goal is to show that, with high probability, GreedySearch converges to an approximate nearest neighbor in $O(\log n)$ steps.
@@ 5. Search Bounds
To analyze the expected path length $k$, we analyze the distribution of neighbors around the destination $v_{k}$, they are uniformly distributed in a $d$-dimensional ball of radius $\rho_{0}$. We partition this ball into concentric annular layers and let $\eta_{i}$ denote the number of neighbors lying in the $i$-th layer. Each $\eta_{i}$ follows a binomial distribution proportional to the shell volume, and the total number of neighbors is bounded by $\eta=O(n^{2/3+\epsilon})$ from Theorem 4.
@@
Using volume arguments, we upper bound the expected total number of neighbors located within the ball of radius $\|v_{k}-v_{0}\|$ by the expected number of dataset points that fall into the same region. Specifically, we show
(20) $\mathbb{E}[k\cdot\eta]\leq\frac{n(n+1)^{(d-1)/d}\rho_{0}^{d+1}\cdot E\left[\left((\rho_{0}-\triangle r)/\rho_{0}\right)^{d(k-1)}\right]}{\triangle r^{d+1}\cdot d},$
where $\triangle r$ is the minimum distinguishable pairwise distance difference among all point triples. Applying Jensen’s inequality (Chernoff), we define an auxiliary function $g(x)$
(21) $g(x):=\eta\cdot x-\frac{n(n+1)^{(d-1)/d}\rho_{0}^{d+1}}{d\cdot\triangle r^{d+1}}\left[-\left(\frac{\rho_{0}-\triangle r}{\rho_{0}}\right)^{d(x-1)}+2\right],$
whose root bounds $\mathbb{E}[k]$. We demonstrate $g(k)<0$, note that $g(x)$ is monotonically increasing, and $g(\log n)>0$, which together imply that $k<\log n$, hence $\mathbb{E}[k]=O(\log n)$. ∎
6. Truncation Parameter Optimization
In the preceding sections, we establish theoretical guarantees for the SNG structure, including sublinear degree growth and logarithmic search complexity. We now leverage these insights to address a practical issue in index construction: how to efficiently select the truncation parameter $R$, which directly controls the trade-off between construction cost and search performance.
6.1. Complexity Analysis
We summarize the full SNG construction pipeline as implemented in practice, incorporating some widely used optimizations (NEURIPS2019; NSG19). The process begins by initializing a randomized $R$-regular graph. For each data point $p$, a GreedySearch procedure is performed from a fixed entry point (e.g., the dataset centroid) to collect an initial candidate set for pruning, with time complexity $T_{\text{search}}=O(R\log n).$ Importantly, in practical implementations the initial candidate set is not the entire dataset; instead, it consists only of the points returned by GreedySearch. This design significantly reduces construction cost compared to the theoretical baseline that assumes full connectivity. In particular, the size of the initial candidate set satisfies $|S_{0}|=O(R\log n)$ in expectation. The candidate set is then refined through iterative SNG pruning, where the truncation parameter $R$ bounds the number of pruning iterations. Let $|S_{0}|$ denote the initial candidate set size; the pruning cost is $T_{\text{prune}}=O(R\cdot|S_{0}|).$ To further enhance graph connectivity and preserve monotonic searchability, reverse edges are inserted after pruning. These insertions may temporarily violate the degree constraint, triggering additional pruning steps. The amortized cost of reverse-edge handling can be modeled as $T_{\text{reverse}}=O(\alpha\cdot R^{3}),$ where $\alpha$ is the pruning parameter controlling sparsification aggressiveness.
The computational cost of each stage—GreedySearch, SNG pruning, and reverse-edge handling—can be analyzed independently. Aggregating these components yields the overall construction complexity, as expressed below.
$n\bigl(C_{1}\cdot R\log n+b_{1}+C_{2}\cdot(R^{2}\log n/\alpha)+b_{2}+C_{3}\cdot(\alpha R^{3})+b_{3}\bigr),$
where $C_{i}$ and $b_{i}$ are implementation-dependent constants.
This expression highlights the competing effects of $R$ and $\alpha$: increasing $R$ improves graph quality but amplifies both pruning and reverse-edge costs.
6.2. Optimization Principle
As discussed in Section 2.3, in practice, $\alpha$ is typically fixed in the range $1\leq\alpha\leq 2$. In contrast, the truncation parameter $R$ introduces an intrinsic trade-off: larger $R$ generally improves recall and robustness of search, but also significantly increases construction cost.
We propose to compute the value of $R$ that achieves optimal efficiency for a given target $\alpha$. We refer to this approach as the marginal optimality principle. Concretely, by differentiating the construction complexity with respect to $\alpha$ and solving for the stationary condition, our theoretical analysis yields the relationship $R^{*}\propto\frac{\log n}{\alpha^{2}}.$ This relation reveals that the optimal truncation parameter scales logarithmically with the dataset size and inversely with the square of the pruning parameter, with the dimensionality $d$ absorbed into the proportionality constant.
@@ 6.2. Optimization Principle
[…]
Input: test pruning parameter $\alpha_{1}$, target pruning parameter $\alpha_{2}$, number of data points $n$
Output: optimized truncation parameter $R^{*}$
[…]
$R\leftarrow n^{2/3}$; // theoretical upper bound
[…]
return $R^{*}$
Algorithm 3 Optimization of Truncation Parameter $R$ of SNG
12
12
12
12
12
12
12
12
12
12
12
12
@@
The goal of Algorithm 3 is to determine the dataset-dependent constant of proportionality in the above relationship. We begin with a reference construction using $R=n^{2/3}$ under a test pruning parameter $\alpha_{1}$, which incurs much lower cost than parameter sweeping. This choice corresponds to the theoretical upper bound on the SNG degree and closely approximates the structure of a non-truncated graph. We then measure the average out-degree $\bar{R}$ of the constructed graph, which captures the intrinsic sparsity induced by the data distribution. Using this estimate, we directly compute the proportionality constant $K^{\prime}$ and the optimized truncation parameter $R^{*}$ for the target pruning parameter $\alpha_{2}$ in closed form.
7. Experiments
We compare a conventional parameter-sweeping pipeline with our analytical parameter determination strategy, and assess both construction efficiency and query performance across different graph-based ANNS algorithms and representative datasets.
7.1. Experimental Setup
7.1.1. Datasets
We evaluate our method on six datasets with diverse dimensionalities and application domains, as summarized in Table 2.
Table 2. Summary of Datasets Used in Experiments
Dataset Cardinality Dimension Query Size Data Type
MSong 992,272 420 200 Audio
DEEP1M 1,000,000 256 1,000 Image
SIFT1M 1,000,000 128 10,000 Image
GIST1M 1,000,000 960 1,000 Image
GloVe 1,193,514 200 1,000 Text
UNIFORM 1,000,000 128 10,000 Synthetic
The first five datasets are standard benchmarks in the ANNS literature and cover a broad range of real-world applications, including image, audio, and text. MSong (msong) consists of audio feature vectors for music similarity search. DEEP1M (DEEP1M) and SIFT1M (DBLP:journals/pami/JegouDS11) are widely used image-retrieval benchmarks based on deep neural features and SIFT descriptors, respectively. GIST1M (DBLP:journals/pami/JegouDS11) contains high-dimensional GIST descriptors. GloVe (glove) provides word embeddings for semantic search.
In addition, we include a synthetic dataset, UNIFORM, where points are generated uniformly at random in a unit hypercube. This dataset is chosen to stress the construction procedure under an idealized and analytically tractable setting. From an information-theoretic perspective, the uniform distribution attains maximum entropy under bounded support and thus exhibits minimal exploitable density variations. Consequently, the pruning dynamics are typically less favorable than in real-world data, where non-uniform density often accelerates effective neighbor elimination. Importantly, our theoretical bounds (Theorem 4) are proved under precisely this distribution, making UNIFORM a controlled setting for empirically verifying that the proposed optimization remains robust even when geometric shortcuts are scarce.
7.1.2. Performance Metrics
We measure both index-construction efficiency and query behavior using several indicators. Index Construction Time measures the end-to-end time to obtain a finalized graph index, which includes both the time spent on selecting construction parameters and the subsequent construction phase.
For query-time evaluation, Recall@10 measures the fraction of queries for which at least one true nearest neighbor appears in the top-10 retrieved results, formally defined as $\text{Recall@}k=|A\cap A^{\prime}|/|A^{\prime}|$, where $A$ is the retrieved set and $A^{\prime}$ is the ground truth with $|A|=|A^{\prime}|=k$. Queries Per Second (QPS) measures throughput as the number of queries processed per second and Query Latency reports the average time per query in milliseconds at different recall levels.
7.1.3. Parameter Settings
For each fixed $\alpha$, we vary the search width $L$ (and analogous search-time controls) to generate recall-latency trade-off. For baseline methods that tune the truncation parameter $R$ via parameter sweeping, we employ a binary-search (or grid-search) strategy guided by a normalized 1:1 weighted loss over query recall and latency, which reflects standard practice in iterative tuning pipelines. In contrast, our approach determines $R$ directly from the analytical relationship derived in Section 6.2.
7.1.4. Computing Environment
All experiments are conducted on a workstation equipped with an Intel Core i7-14700KF CPU (20 cores, 3.4GHz base and up to 5.6GHz boost), 64GB DDR4-3200 memory, and a Samsung 990 EVO Plus 2TB NVMe SSD, running Ubuntu 22.04 LTS. We compile all code with GCC 11.4 and enable -O3 optimizations.
7.1.5. Compared Algorithms
We evaluate three representative SNG-based ANNS algorithms and apply our optimization principle consistently across them. Vamana (NEURIPS2019), which underlies DiskANN-style indexing, relies on SNG-like pruning and exposes the parameters $\alpha$ and $R$ that directly control graph sparsification and maximum out-degree. NSG (NSG19) constructs a navigable graph with a bounded out-degree constraint and typically requires careful tuning of $R$ to balance construction cost and search quality. HNSW (HNSW2020) builds a hierarchical multi-layer graph, where the construction-time candidate list size (e.g., efConstruction) plays a similar role in controlling construction overhead and the resulting graph connectivity. We denote the variants using our proposed parameter optimization strategy as OPT-DiskANN, OPT-NSG, and OPT-HNSW, and refer to their original implementations as DiskANN, NSG, and HNSW, respectively.
7.2. Index Construction Performance
We begin by comparing the end-to-end index construction time, including both parameter tuning and the subsequent graph building. Figure 8 reports results across all datasets, where the stacked bars separate tuning time from construction time. Across the board, the most salient difference is the tuning component: parameter sweeping repeatedly triggers expensive partial or full constructions to probe candidate $R$ values, whereas our method largely removes this iterative overhead by computing $R$ analytically once the necessary reference statistics are obtained. As a result, the total construction time decreases substantially even when the actual build phase remains comparable.
Table 3. Index Construction Speedup of OPT-SNG over Baseline Methods
Dataset Speedup over Baseline Avg
Vamana NSG HNSW
[…]
Table 3 summarizes the speedups achieved over binary-search baselines. We observe consistent improvements across datasets and algorithms, ranging from $2.2\times$ to $15.4\times$, with an average speedup of $5.9\times$. The gains are particularly pronounced on GloVe , where NSG achieves a $15.4\times$ speedup, and on SIFT1M, where Vamana achieves $7.9\times$. These large gains occur because the binary-search pipeline typically requires more iterations to converge to a good truncation level on these datasets, leading to large cumulative tuning costs.
7.3. Query Performance Evaluation
We next evaluate whether the analytically chosen parameters (OPT) preserve query-time behavior compared to the parameters selected by parameter sweeping. We report recall and QPS trends by varying the search width $L$ under each $\alpha$.
Table 4. Recall@10 comparison at fixed latency targets. Latency thresholds are dataset- and algorithm-specific, chosen to reflect typical operating points.
Dataset Vamana NSG HNSW
Binary OPT Binary OPT Binary OPT
MSong 99.26 99.32 99.79 99.82 86.27 91.85
DEEP1M 96.77 98.76 97.94 99.33 98.93 99.71
SIFT1M 99.84 99.90 99.64 99.70 99.92 99.99
GIST1M 99.63 99.76 96.16 96.79 98.20 99.63
GloVe 88.91 89.94 90.34 91.83 92.97 94.61
UNIFORM 98.78 99.62 93.25 95.96 96.53 97.64
@@ 7.3. Query Performance Evaluation
Figure 3. Query performance across varying $\alpha$ values on real-world datasets
Figures 3 present recall versus QPS curves under different $\alpha$ settings. Across datasets, the curves produced by OPT closely track and in several cases slightly improve upon the curves of the original baseline, indicating that removing iterative tuning does not come at the expense of query quality. Intuitively, this behavior is consistent with our derivation: the analytical rule targets the better construction-quality regime that the parameter sweep attempts to locate empirically, but avoids repeatedly paying the cost of probing suboptimal intermediate choices.
To provide a compact comparison under fixed latency budgets, Table 4 summarizes the achieved Recall at representative, dataset and algorithm specific latency targets. Specifically, the latency budgets are set as follows: Vamana—5ms (DEEP1M), 200ms (GIST1M, GloVe, MSong), and 10ms (UNIFORM, SIFT1M); NSG—5ms (MSong), 10ms (GloVe, UNIFORM), 0.15ms (SIFT1M), 5ms (GIST1M), and 2ms (DEEP1M); HNSW—0.5ms (SIFT1M), 5ms (GIST1M), 1ms (DEEP1M), 10ms (UNIFORM), 15ms (GloVe), and 10ms (MSong). Bold entries indicate the better result between Binary Search and OPT for each algorithm–dataset pair.
The results show that OPT consistently matches or improves Recall@10 across all three algorithms and all datasets. The improvements are most pronounced on DEEP1M, MSong, GIST1M, and GloVe, where analytical parameter selection avoids locally optimal but globally suboptimal truncation choices. Notably, HNSW exhibits the largest recall gains on SIFT1M, while on the challenging UNIFORM dataset, the most significant improvement is observed for NSG.
7.4. Discussion on UNIFORM: A Worst-Case Construction Regime
As shown in Table 2, UNIFORM shares the same scale and dimensionality as SIFT1M, enabling a controlled comparison that isolates the effect of data distribution.
Figure 4 shows that the UNIFORM dataset consistently yields lower throughput than SIFT1M across all evaluated algorithms at comparable recall levels. For example, at 99% Recall@10, Vamana on SIFT1M achieves approximately $7\times$ higher QPS than on UNIFORM. Even under this challenging distribution, OPT consistently achieves higher QPS than the baseline at the same recall levels, demonstrating that our optimization remains effective and preserves query efficiency under uniform data without favorable geometric structure.
Figure 8 further demonstrates that UNIFORM produces systematically larger average node degrees than SIFT1M. This behavior precisely characterizes a worst-case regime for SNG construction. Our theoretical degree bound of $O(n^{2/3+\epsilon})$ (Theorem 4) is derived under such distributions, since the uniform distribution represents the least favorable case, this bound extends naturally to other distributions with more structured density. Empirically, even without truncation, the maximum degree observed on UNIFORM is 3873, which remains well within the theoretical bound.
These observations are consistent with our theoretical analysis. Uniformly distributed points lack the intrinsic clustering structure commonly present in real-world datasets, which reduces opportunities for aggressive pruning and geometric shortcutting. As a result, the graph construction must introduce more edges to capture neighborhood relationships, leading to increased pruning iterations, higher node degrees, and reduced search efficiency.